OpenStudy (anonymous):

algebra 1- system of equations 1.Tony and Belinda have a combined age of 56. Belinda is 8 more than twice Tony’s age. How old is each?

OpenStudy (anonymous):

First solve for x to find tony's age. 56= 8+2x 2x= 56-8 2x=48 x=24 (tony's age is 24) to find belinda's age, just add 8 to 24. 24+8= 32 ANSWER: Tony is 24 Belinda is 32

OpenStudy (anonymous):

t@ kldcm ur solution is nt right

OpenStudy (anonymous):

let tony age=x belinda age=y then x+y=56 and y=2x+8 by solving this we get x=16 y=40

OpenStudy (anonymous):

ah yes! MR Crazzy is right! i misread the question!